Consider a boy on a trolley who throws a ball with speed 20m/s at an angle 37 degrees w.r.t trolley in direction of motion of trolley which moves horizontally with speed 10m/s, find range w.r.t ground.

Grade:

1 Answers

Sumit Majumdar IIT Delhi
9 years ago
Dear student,
The net horizontal velocity would be given by:
u_{x}=20cos37^{\circ}+10; u_{y}=20sin37^{\circ}
Hence, range would be given by:
R=\frac{\left ( 20cos37^{\circ}+10 \right )\left ( 2 \times 20sin37^{\circ}\right )}{g}=\frac{40}{g}sin37^{\circ}\left ( 20cos37^{\circ}+10 \right )=62.5 mRegards
Sumit